What is the molar concentration of a solution of 63.0g HNO3 in 0.25 L of H2O?
1
Expert's answer
2020-11-22T11:11:11-0500

find the n:

"n=\frac{m}{M}=\frac{63}{63.012}=0.99"


 the molar concentration of a solution

"c=\frac{n}{V}=\frac{0.99}{0.25}=3.96"